• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 16:09:15 on August 8,2005

    Entry number 151164

    keyword=Day Shift Run Summary

    8:00 Shift started with Right Arm Q2 down. 
    
    08:46 Ed Returned from the Hall. He says there is a temperature and 
    communication problem, he wants to bring Mark in on the problem. 
    
    09:18 Ed did a hard reboot of the IOC. We are cycling the quads and preparing 
    to go back to production. 
    
    09:29 Ed is done with work in the Hall. We ask for Beam.
    
    09:31 We lost the Right Arm IOC again.
    
    09:32 Reboot the IOC.
    
    09:43 Start taking beam on the Left Arm only.
    
    10:08 Start Compton Run
    
    Left arm only run
    4214   09:44-10:14 
    
    Production runs
    4215   10:15-11:10    100k
    4217   11:53-12:49    100k
    4218   12:49-13:31     75k
    4220   13:51-14:46    100k
    4221   14:47-15:43    100k
    4222   15:44-
    
    Right arm only run
    4219  13:32-13:50      33k
    
    
    11:13 Request an access to power cycle the Compton crate. 
    
    11:43 We power cycled the network switch. Sirish says that the Compton looks ok.
    
    11:53 Beam returns
    
    12:10 target problems, reboot IOC,
    
    12:18 Beam returns
    
    13:06 Rebot Target IOC
    
    13:30 Left arm ioc crashed, reboot.
    
    13:50 Left Arm back up.
    
    15:50 Compton ROCs not responding again.
    
    16:00 shift ended running production of the last IHWP-OUT run.